How hot does a sample need to be to have the given percentage of its molecules in the first excited vibrational state? Use as the separation between the ground and first excited energy levels the value given in the question. [ ε=1.3199×10^3 cm−1, X=3.7201×10−1K.]
Answer #1

Numerical analysis of ro-vibrational spectral data would appear to be complicated by the fact that the wavenumber for each transition depends on two rotational constants, {\displaystyle B^{\prime \prime }} and {\displaystyle B^{\prime }}. However combinations which depend on only one rotational constant are found by subtracting wavenumbers of pairs of lines (one in the P-branch and one in the R-branch) which have either the same lower level or the same upper level.[2][3] For example, in a diatomic molecule the line denoted P(J + 1) is due to the transition (v = 0, J + 1) → (v = 1, J), and the line R(J − 1) is due to the transition (v = 0, J − 1) → (v = 1, J). The difference between the two wavenumbers corresponds to the energy difference between the (J + 1) and (J − 1) levels of the lower vibrational state and denoted by {\displaystyle \Delta _{2}} since it is the difference between levels differing by two units of J. If centrifugal distortion is included, it is given by[4]

{\displaystyle \Delta _{2}^{\prime \prime }F(J)={\bar {\nu }}[R(J-1)]-{\bar {\nu }}[P(J+1)]=(2B^{\prime \prime }-3D^{\prime \prime })\left(2J+1\right)-D^{\prime \prime }\left(2J+1\right)^{3}}

The rotational constant of the ground vibrational state B′′ and centrifugal distortion constant, D′′ can be found by least-squares fitting this difference as a function of J. The constant B′′ is used to determine the internuclear distance in the ground state as in pure rotational spectroscopy. (See Appendix)

Similarly the difference R(J) − P(J) depends only on the constants B′ and D′ for the excited vibrational state (v = 1), and B′ can be used to determine the internuclear distance in that state (which is inaccessible to pure rotational spectroscopy).

{\displaystyle \Delta _{2}^{\prime }F(J)={\bar {\nu }}[R(J)]-{\bar {\nu }}[P(J)]=(2B^{\prime }-3D^{\prime })\left(2J+1\right)-D^{\prime }\left(2J+1\right)^{3}}